In a local market, the monthly price of Internet access service decreases from $20 to $10, and the total quantity of monthly accounts across all Internet access providers increases from 90,000 to 190,000 What is the value price elasticity of demand, expressed as a positive number?(Round your answer to two decimal places)
Answer #1

Price elasticity of demand = % change in quantity / %change in price

% change in quantity = (190000-90000)/90000 = 100000/90000 = 10/9

% Change in Price = (10-20)/20 = -10/20 = -0.5

Price Elasticity of demand = 10/9*(-0.5) = 10/4.5 = -2.22

Here we will ignore the negative sign.

Price Elasticity of Demand = 2.22